OpenStudy (anonymous):

Find the standard form of the equation of the parabola with a focus at (0, -10) and a directrix at y = 10

OpenStudy (jdoe0001):

keeping in mind that the distance from the vertex to the focus, is the exact same one as the distance from the vertex to the directrix, in the opposite direction  where do you think is the vertex?

OpenStudy (anonymous):

is it 10?

OpenStudy (jdoe0001):

well, if the distance from the vertex to the focus, is the exact same one as the distance from the vertex to the directrix that means that the vertex is half-way between both of them

OpenStudy (anonymous):

So 5

OpenStudy (jdoe0001):



OpenStudy (jdoe0001):

see now where the vertex is?

OpenStudy (anonymous):

0?

OpenStudy (jdoe0001):

yeap, is at the origin, 0, 0 notice the distance from the vertex to the focus (0, -10), is 10 units the "focus form" for a parabola like so is \(\bf (x-h)^2=4p(y-k)\) (h, k) = vertex coordinates p = distance from the vertex to the focus point

OpenStudy (anonymous):

so i just plug the numbers in? (0 - 0)^2 = 4(10) (-10 - 0)

OpenStudy (jdoe0001):

well, \(\bf (x-h)^2=4p(y-k) \implies (x-0)^2=4(10)(y-0)\)

OpenStudy (jdoe0001):

yes, and then you can solve for "y" or equate to 0

OpenStudy (jdoe0001):

well, it has an "x" and "y", so might as well solve for "y"

OpenStudy (anonymous):

x^2 - 40 = y

OpenStudy (jdoe0001):

well, \(\bf \bf (x-h)^2=4p(y-k) \implies (x-0)^2=4(10)(y-0) \implies x^2 = 40y\\ \cfrac{x^2}{40} = y \implies \cfrac{1}{40}x^2 = y \)

OpenStudy (jdoe0001):

well, that's the parabola equation

OpenStudy (anonymous):

Oh okay. thank you!

OpenStudy (jdoe0001):

yw